OpenStudy (anonymous):

Give the sum of 2x -3 and 7x + 21

OpenStudy (anonymous):

\[2x+7x=?\]

OpenStudy (anonymous):

9x...

OpenStudy (anonymous):



OpenStudy (adryin):

Well 2x-3 and 7x+21 add up to 180 degrees.

OpenStudy (adryin):

(2x+7x)+(21-3) = 180

OpenStudy (adryin):

9x+18 = 180

OpenStudy (adryin):

9x = 162

OpenStudy (adryin):

x = 18

OpenStudy (anonymous):

Thank you so much!